OpenStudy (anonymous):

find the slope of (3,11) and (-1,12)

hero (hero):

Do you know the slope formula?

OpenStudy (anonymous):

Slope of the line

hero (hero):

The slope formula is \(m = \dfrac{y_2 - y_1}{x_2 - x_1}\)

OpenStudy (anonymous):

I got this

OpenStudy (anonymous):

Gimme a sec

OpenStudy (anonymous):

Crap, guess he got it first. If you want to find the slope of those coordinates just plug em into that equation hero just posted. Simple as that.

hero (hero):

So the first point is \((x_1, y_1) = (3,11)\) The second point is \((x_2, y_2) = (-1,12)\) When you plug those values in to the slope formula you get: \(m = \dfrac{12 - 11}{-1 - 3}\) After simplification you get \(m = -\dfrac{1}{4}\) So the slope of the line is \(-\dfrac{1}{4}\)

hero (hero):

@squintana3